### v4

# Bejing 100048, People’s Republic of China

Abstract
It was shown by Kutnar and Šparl in 2009 that every connected vertex-transitive graph of order $6p$, where $p$ is a prime, contains a Hamilton path. In this paper, it will be shown that every such graph contains a Hamilton cycle, except for the triangle-replaced graph of the Petersen graph.
1 Introduction
A simple path (reps. cycle) going through all vertices of a graph is called a Hamilton path (resp. cycle). In 1969, Lovász [22] asked that
is there any finite connected vertex-transitive graph of order more than 3 without a Hamilton path ?
Till now, no such graph is known to exist. With the exception of $K_{2}$, only four connected vertex-transitive graphs that do not have a Hamilton cycle are known to exist. These four graphs are the Petersen graph, the Coxeter graph and the two graphs obtained from them by replacing each vertex by a triangle. The fact that none of these four graphs is a Cayley graph has led to a folklore conjecture that except for $K_{2}$, every Cayley graph has a Hamilton cycle, see [4, 8, 9, 11, 12, 16, 24, 34, 35].
The existence of Hamilton paths, and in some cases also Hamilton cycles, in connected vertex-transitive graphs has been shown for graphs of particular orders, such as, $kp$, where $k\leq 6$, $p^{j}$, where $j\leq 5$, and $2p^{2}$, where throughout this paper $p$ denotes a prime, see [3, 6, 7, 17, 18, 19, 25, 26, 28, 29, 30, 31, 32, 33, 36] and a survey paper [16].
Recently, Kutnar, Marušič and the first author proved that every connected vertex-transitive graph of order $pq$, where $p$ and $q$ are primes, has a Hamilton cycle, except for the Petersen graph (see [6]). As for the connected vertex-transitive graphs of order $2pq$, where $p$ and $q$ are primes, Tian, Yu and the first author showed that except for the Coxeter graph, every graph of such order contains a Hamilton cycle, provided the automorphism group acts primitively on its vertices (see [7]).
In [19], Kutnar and Šparl showed that every connected vertex-transitive graph of order $6p$, where $p$ is a prime, contains a Hamilton path, and in some particular cases also contains a Hamilton cycle. The main result of this paper is the following theorem.
Theorem 1.1
Except for the triangle-replaced graph of the Petersen graph, every connected vertex-transitive graph of order $6p$ contains a Hamilton cycle, where $p$ is a prime.
This paper is organized as follows: after this introductory section, some preliminaries will be given in Section 2 and the proof of Theorem 1.1 will be given in Section 3.
2 Terminology, notation and some useful results
This section consists of four subsections dealing with basic definitions and notation, generalized orbital graphs, existence of Hamilton cycles in particular graphs, and group- and finite field-theoretic results, respectively.
2.1 Basic definitions and notation
Throughout this paper, graphs are finite, simple, connected and undirected, and groups are finite. Furthermore, a multigraph is a generalization of a graph in which multiple edges and loops are allowed. Given a graph $X$, we denote the vertex set by $V(X)$ and the edge set by $E(X)$, respectively. For two adjacent vertices $u,v\in V(X)$, we write $u\sim v$ and denote the corresponding edge by $\{u,v\}$. Let $U$ and $W$ be two disjoint subsets of $V(X)$. Then by $X(U)$ and $X[U,W]$ we denote the subgraph of $X$ induced by $U$ and the bipartite subgraph of $X$ induced by the edges having one end-vertex in $U$ and the other end-vertex in $W$, respectively.
Given a group $G$ acting faithfully and transitively on a set $V$. A nonempty subset $B$ of $V$ is called a $block$ for $G$ if for each $g\in G$ either $B^{g}=B$ or $B^{g}\cap B=\emptyset$, where $V$ and the singletons $\{v\}$ $(v\in V)$ are called the trivial blocks, and other blocks are called nontrivial blocks. Put =$\mathcal{B}=\{B^{g}:g\in G\}$. Then the sets in $\mathcal{B}$ form a partition of $V$. We call $\mathcal{B}$ the system of blocks containing $B$. If $G$ has no nontrivial blocks on $V$, then $G$ is primitive, and is imprimitive otherwise. If every nontrivial normal subgroup of $G$ is transitive on $V$, then $G$ is quasiprimitive, and is non-quasiprimitive otherwise. If $G$ acts non-quasiprimitively on $V$ with an intransitive normal subgroup $N$, then for any $v\in V$, $v^{N}=\{v^{g}:g\in N\}$ is called an $N$-block. The set of $N$-blocks forms a system of blocks for $G$.
A graph $X$ is said to be vertex-transitive if its automorphism group Aut ($\hbox{\rm Aut\,}(X)$ acts transitively on $V(X)$. A vertex-transitive graph $X$ for which every transitive subgroup of Aut ($\hbox{\rm Aut\,}(X)$ is primitive is called a primitive graph and an imprimitive graph otherwise. If $X$ is imprimitive with an imprimitivity block system induced by a normal subgroup of some transitive subgroup $G\leq\hbox{\rm Aut\,}(X)$, then the graph $X$ is said to be genuinely imprimitive. If $X$ is imprimitive, but there exists no transitive subgroup $G$ which having an intransitive normal subgroup, then $X$ is said to be quasiprimitive.
Let $m\geq 1$ and $n\geq 2$ be integers. An automorphism $\rho$ of a graph $X$ of order $n$ is called $(m,n)$-semiregular (in short, semiregular) if as a permutation on $V(X)$ it has a cycle decomposition consisting of $m$ cycles of length $n$. The question whether all vertex-transitive graphs admit a semiregular automorphism is one of the famous open problems in algebraic graph theory (see, for example, [1, 2, 5, 10, 27]). Let $\mathcal{P}$ be the set of orbits of $\langle\rho\rangle$. Let $A,B\in\mathcal{P}$. By $d(A)$ and $d(A,B)$ we denote the valency of $X(A)$ and $X[A,B]$, respectively, reminding that the graph $X[A,B]$ is regular. Let the quotient graph corresponding to $\mathcal{P}$ be the graph $X_{\mathcal{P}}$ whose vertex set is $\mathcal{P}$ with $A,B\in\mathcal{P}$ adjacent if there exist vertices $a\in A$ and $b\in B$, such that $a\sim b$ in $X$. Let the quotient multigraph corresponding to $\rho$ be the multigraph $X_{\rho}$ whose vertex set is $\mathcal{P}$ and in which $A,B\in\mathcal{P}$ are joined by $d(A,B)$ edges. Note that the quotient graph $X_{\mathcal{P}}$ is precisely the underlying graph of $X_{\rho}$.
A graph $X$ is called an $(m,n)$-metacirculant, where $m$ and $n$ are positive integers, if $X$ is of order $|V|=mn$ and has two automorphisms $\rho$, $\sigma$ such that
(1) $\langle\rho\rangle$ has the number $m$ of semiregular orbits on $V$;
(2) $\sigma$ cyclically permutes the number $m$ of orbits of $\langle\rho\rangle$ and normalizes $\langle\rho\rangle$; and
(3) $\sigma^{m}$ fixes at least one vertex of $X$.
2.2 Generalized orbital graphs
A permutation group $G$ on a set $V$ induces the action of $G$ on $V\times V$, whose orbits are called orbitals of $G$, while $O_{0}=\{(x,x)\colon x\in V\}$ is said to be trivial. Every orbital $O_{i}$ corresponds to a paired orbital $O_{i}^{*}=\{(x,y):(y,x)\in O_{i}\}$, and $O_{i}$ is said to be self-paired if $O_{i}=O_{i}^{*}$.
For an arbitrary union $\cal{O}$ of orbitals (having empty intersection with $O_{0}$), the generalized orbital (di)graph $X(V,\cal{O})$ of the action of $G$ on $V$ with respect to $\cal{O}$ is a simple (di)graph with the vertex set $V$ and edge set $\cal{O}$. It can be viewed as an undirected graph if and only if $\cal{O}$ coincides with its symmetric closure, that is, $\cal{O}$ has the property that $(x,y)\in\cal{O}$ implies $(y,x)\in\cal{O}$. Further, $X(V,\cal{O})$ is said to be basic orbital graph if ${\cal O}=O_{i}\cup O_{i}^{*}$ for some $O_{i}$.
Let $X$ be a graph with a transitive group $G$ of Aut ($\hbox{\rm Aut\,}(X)$. By $G_{v}$ and $[G:H]$ we denote the stabilizer of $v\in V(X)$ and the set of cosets of $G$ relative to $H$, respectively, where $H=G_{v}$. Then there exists a 1-1 correspondence between orbitals and suborbits of $G$, that is orbits of $H$ on $V(X)$. A suborbit corresponding to a self-paired orbital is said to be self-paired. Therefore, the (generalized) orbital (di)graph $X(G/H,\cal{O})$ is denoted by $X(G,H,\mathcal{S})$ too, provided $\cal{O}$ corresponds to a union $\mathcal{S}$ of suborbits, while $\mathcal{S}$ may be replaced by a union $D$ of some double cosets of $H$. In $X(G,H,D)$, $Hg_{1}\sim Hg_{2}$ if and only if $Hg_{2}\in\{Hdg_{1}:d\in D\}$, where $g_{1},g_{2}\in G$.
Let $X$ be a genuinely imprimitive graph of order $n$ and let $G$ be an imprimitive subgroup of Aut ($\hbox{\rm Aut\,}(X)$ with a nontrivial intransitive normal $N$. Then $N$ induces blocks of length $m$, where $m\geq 2$ and $m\mid n$. Then $X\cong X(G,H,D)$ for the union of some double cosets of $H$, where $H=G_{v}$. Let $\mathcal{B}$ be the set of $N$-blocks. Let $A,B\in\cal{B}$. Let the quotient graph corresponding to $\cal{B}$ be the graph $X_{\mathcal{B}}$ whose vertex set is $\cal{B}$ with $A,B\in\cal{B}$ adjacent if there exist vertices $a\in A$ and $b\in B$, such that $a\sim b$ in $X$. $X_{\mathcal{B}}$ is also called an $m$-block graph induced by $N$. Let $\overline{G}=G/N$, $\overline{H}=HN/N$ and $\overline{D}=\{\overline{HdH}:d\in D\}$. Then $X_{\mathcal{B}}\cong X(\overline{G},\overline{H},\overline{D})$ implying that $X_{\mathcal{B}}$ is a vertex-transitive graph. Moreover, $X_{\mathcal{B}}$ is connected if $X$ is connected.
2.3 Existence of Hamilton cycles in particular graphs
The following known results about existence of Hamilton cycles in particular graphs will be used later.
Proposition 2.1
[15, Theorem 1] Every $2$-connected regular graph of order $n$ and valency at least $n/3$ contains a Hamilton cycle.
Proposition 2.2
[19, Lemma 2.4] Let $X$ be a $G$-imprimitive graph of order $mq$, where $q$ is a prime, and let $G$ be an imprimitive subgroup of Aut ($\hbox{\rm Aut\,}(X)$ and $N$ be a normal subgroup of $G$ with orbits of length $q$. Then $X$ has an $(m,q)$-$semiregular$ automorphism whose orbits coincide with the orbits of $N$.
Lemma 2.3
With the same notation as in Proposition 2.2, suppose that $G/N$ is faithful on the set of $N$-blocks. Let $N_{0}=\langle N,P\rangle$ and $G_{0}=\langle G,P\rangle$, where $P$ is the set of all $(m,q)$-$semiregular$ automorphisms whose orbits coincide with the orbits of $N$ in Aut ($\hbox{\rm Aut\,}(X)$. Then $N_{0}\lhd G_{0}$ and $G/N\cong G_{0}/N_{0}$.
Proof   Let $\mathcal{B}$ be the set of $N$-blocks. Then for every $g\in G$ and $\rho\in P$, we have $B^{\rho^{g}}\in\mathcal{B}$, where $B\in\mathcal{B}$. Set $fix(\rho^{g})$, the set of fixed points of $\rho^{g}$ in $V(X)$. Then $fix(\rho^{g})=fix(\rho)=\emptyset$ and so $\rho^{g}$ is an $(m,q)$-semiregular automorphism whose orbits coincide with the orbits of $N$. This implies that $\rho^{g}\in P$ and $P^{g}=P$. Consequently, $\langle N^{g},P^{g}\rangle=\langle N,P\rangle$ and $N_{0}\lhd G_{0}$. Moreover, we have $G_{0}/N_{0}=GN_{0}/N_{0}\cong G/(N_{0}\cap G)=G/N$. $\Box$
By Lemma 2.3, if $X$ is a $G$-imprimitive graph as described above, we always assume that $G$ contains an $(m,q)$-$semiregular$ automorphism $\rho$ and $\rho\in N$.
Proposition 2.4
[29, Lemma 5] Let $X$ be a graph admitting an $(m,p)$-$semiregular$ automorphism $\rho$, where $p$ is a prime. Let $C$ be a cycle of length $k$ in the quotient graph $X_{\mathcal{P}}$, where $\mathcal{P}$ is the set of orbits of $\rho$. Then the lift of $C$ either contains a cycle of length $kp$ or it consists of $p$ disjoint $k$-cycles. In the latter case, we have $d(S,S^{{}^{\prime}}$) = 1 for every edge $SS^{{}^{\prime}}$ of $C$.
Proposition 2.5
[23, Theorem 1.2] Let $G$ be a finite group. Every connected Cayley graph on $G$ has a Hamiltonian cycle if $G$ is of order
$kp,1\leq k\leq 47;\quad kpq,1\leq k\leq 7;\quad pqr;\quad kp^{2},1\leq k\leq 4;$
$kp^{3},\,1\leq k\leq 2;\quad p^{k},1\leq k\textless\infty,$
where $p$, $q$ and $r$ are distinct prime.
Proposition 2.6
[6, Theorem 1.4] With the exception of the Petersen graph, a connected vertex-transitive graph of order $pq$, which $p$ and $q$ are primes, contains a Hamilton cycle.
Proposition 2.7
[19, Theorem 1.1] Every connected vertex-transitive graph of order $6p$, where $p$ is a prime, contains a Hamilton path. Moreover, with the exception of the truncation of the Petersen graph, every such graph which is not genuinely imprimitive contains a Hamilton cycle.
Proposition 2.8
[7, Theorem 1.1] Except for the Coxeter graph, every connected vertex-transitive graph of order $2rs$ contains a Hamilton cycle provided the automorphism group acts primitively on its vertices, where $r$ and $s$ are primes.
Proposition 2.9
[19, Lemma 3.4, 3.5] Suppose that $X$ is a connected vertex-transitive graph of order $6p$, where $p$ is a prime. If a transitive subgroup $G$ of $X$ contains a normal subgroup $N$ which induces blocks of length $6$ or $p$, then $X$ has a Hamilton cycle.
Lemma 2.10
[…]
Proof   Let $C:\,H,Hg_{1},Hg_{2},\ldots,Hg_{t-1},H$ be a Hamilton cycle of $X$.
(1) $Y_{1}=X(G,H,Dc\cup Dc^{-1})$.
Since $Hg_{i}c^{j}\sim Hg_{i+1}c^{j+1}$ in $Y_{1}$ and $(t,p)=1$, we may get a Hamilton cycle of $Y_{1}$:
$\begin{array}[]{ll}C_{1}:&H,Hg_{1}c,Hg_{2}c^{2},\ldots,Hg_{t-1}c^{t-1},Hc^{t},Hg_{1}c^{t+1},\ldots,Hg_{t-1}c^{2t-1},\\ &Hc^{2t},Hg_{1}c^{2t+1},\ldots,Hg_{1}c^{(p-1)t+1},\ldots,Hg_{t-1}c^{pt-1},H.\end{array}$
(2) $Y_{2}=X(G,H,D\cup HcH\cup Hc^{-1}H)$. Note that there exists the number $p$ of disjoint cycles of length $t$ in $Y_{2}$:
$Hc^{i},Hg_{1}c^{i},Hg_{2}c^{i},\ldots,Hg_{t-1}c^{i},Hc^{i},$
where $i\in\mathbb{Z}_{p}$. Observing $Hg_{s}c^{i}\sim Hg_{s}c^{i\pm 1}$, we get a Hamilton cycle of $Y_{2}$ as follows:
[…]
Now, suppose that $T$ is primitive. Then the $H$-double coset of $G$ has the form $Hg^{\prime}c^{i}H$ for some $g^{\prime}\in T$ and $i\in\mathbb{Z}_{p}$. Clearly, $\langle H,g^{\prime}c^{i}\rangle=G$ if $g^{\prime}\notin H$ and $i\neq 0$. Suppose that $D^{\prime}=Hg^{\prime}H\cup Hg^{\prime-1}H$ and $D_{0}=Hg^{\prime}c^{i}H\cup Hc^{-i}g^{\prime-1}H$. Firstly, suppose that $o(c)=2$. Then $Hg^{\prime}cH=Hg^{\prime-1}cH$ and so $D_{0}=D^{\prime}c\cup D^{\prime}c^{-1}$. Secondly, suppose that $o(c)\neq 2$. Then we may assume $T$ is of degree $2r$. Since $T$ is primitive, we know that every suborbits of $T$ is self-paired and $D_{0}=D^{\prime}c^{i}\cup D^{\prime}c^{-i}$. Thus, if $X(G,H,D_{0})$ doesn’t contain a subgraph with the form $X(T,H,D^{\prime}c^{k}\cup D^{\prime}c^{-k}))$, then it contains a connected subgraph with the form $X(G,H,D^{\prime}\cup Hc^{k}H\cup Hc^{-k}H)$. $\Box$
2.4 Group- and finite field-theoretic results
Proposition 2.11
[13, Theorem 1.49] Every transitive group $G$ of prime degree $p$ is prright-normal-factor-semidirect-$\mathbb{Z}_{p}\rtimes\mathbb{Z}_{r}$ which is isomorphic to a subgroup of $\hbox{\rm AGL}(1,p)$, where $r\bigm{|}(p-1)$; $\hbox{\rm Soc}(G)=A_{p}$; $\hbox{\rm Soc}(G)=\hbox{\rm PSL}(2,11)$ of degree 11; $\hbox{\rm Soc}(G)=M_{11}$ or $M_{23}$ of degree $11$ or $23$, respectively; or $\hbox{\rm Soc}(G)=\hbox{\rm PSL}(d,q)$ of degree $p=(q^{d}-1)/(q-1)$.
By Proposition 2.11, the following result can be extracted from [32, Proposition 2.5].
Proposition 2.12
Let $p,q$ be primes and $p\textgreater q$ and let $G$ be a quasiprimitive and imprimitive group of degree $pq$. Then $G$ is an almost simple group and $\hbox{\rm Soc}(G)\cong T$, where $T\in\{A_{p},\hbox{\rm PSL}(2,11),M_{11},M_{23},\hbox{\rm PSL}(d,q)\}$, where $d$ is a prime, $(d,q-1)=1$ and $p=(q^{d}-1)/(q-1)$.
The following two results can be extracted from [21].
Proposition 2.13
Every primitive transitive group $G$ of degree $2p$, where $p$ is a prime, has the socle: $A_{5}$ of degree 10; $A_{5}$ of degree 6; $G=M_{22}$ of degree 22; $\hbox{\rm PSL}(2,q)$ of degree $q+1$; or $A_{2p}$ of degree $2p$. Moreover, all of them are $2$-transitive, except for $\hbox{\rm Soc}(G)=A_{5}$ of degree 10.
Proposition 2.14
Every primitive transitive group $G$ of degree $3p$, where $p$ is a prime, has the socle: $A_{3p}$ of degree $3p$; $A_{6}$ of degree $15$; $A_{7}$ of degree $15$; $A_{7}$ of degree 21; $\hbox{\rm PSL}(2,7)$ of degree 21; $\hbox{\rm PSL}(2,19)$ of degree $57$; $\hbox{\rm PSL}(d,q)$ of degree $3p=\frac{q^{d}-1}{q-1}$. Moreover, except for $\hbox{\rm Soc}(G)=A_{6}$, every primitive group of degree $3p$ contains a metacyclic transitive subgroup $T$, where $T\cong\mathbb{Z}_{p}\rtimes\mathbb{Z}_{3^{e}}\lessapprox\hbox{\rm AGL}(1,p)$ or $T\lessapprox\mathbb{Z}_{3p}$.
By Proposition 2.12, the following result can be extracted from [32] and [33, Theorem 2.1].
Proposition 2.15
Let $p$ and $q$ be distinct primes, where $p\textgreater q$, and let $X$ is a connected vertex-transitive graph of order $pq$ with a transitive group $G$, where $G$ is quasiprimitive but not primitive on $V(X)$. Then either Aut ($\hbox{\rm Aut\,}(X)$ contains a metacirculant transitive subgroup which is isomorphic to $\mathbb{Z}_{p}\rtimes\mathbb{Z}_{q^{t}}$; or $G$ contains a transitive subgroup with the socle isomorphic to $\hbox{\rm SL}(2,2^{2^{s}})$, where $p=1+2^{2^{s}}$ is a Fermat prime and $q|(p-2)$, acting on cosets of $\mathbb{Z}_{2}^{2^{s}}\rtimes\mathbb{Z}_{(2^{2^{s}}-1)/q}$.
Proposition 2.16
[14, I. Theorem 7.8] Suppose that $G$ is a finite group with a normal subgroup $H$ and $P$ is a $Sylow$ $p$-subgroup of $H$, then $G=N_{G}(P)H$.
Proposition 2.17
[14, I. Theorem 4.5] For a subgroup $H$ of group $G$, the factor group $N_{G}(H)/C_{G}(H)$ is isomorphic to a subgroup of Aut ($\hbox{\rm Aut\,}(H)$, the group of automorphisms of $H$.
Proposition 2.18
[20, page 331] Let $f\in\mathbb{F}_{q}[x,y]$ be absolutely irreducible and of degree $d$, and let $N$ be the number of solutions of $f(x,y)=0$ in $\mathbb{F}_{q}^{2}$. Then
$|N-q|\leq(d-1)(d-2)q^{\frac{1}{2}}+d^{2}.$
3 The proof of Theorem 1.1
Outline of the proof: To prove Theorem 1.1, let $X$ be a connected vertex-transitive graph of order $6p$, where $p$ is a prime. Then we need to consider the following two cases, separately:
  1. (1)
    Aut ($\hbox{\rm Aut\,}(X)$ contains a transitive subgroup $G$ which contains a normal subgroup $N\neq 1$ inducing nontrivial blocks $B$ on $V(X)$.
  2. (2)
    Every transitive subgroup $G$ of Aut ($\hbox{\rm Aut\,}(X)$ acts quasiprimitively on $V(X)$.
By Proposition 2.7, $X$ contains a Hamilton cycle provided $X$ is in Case (2) except for the triangle-replaced graph of the Petersen graph. If $p=2$ or $3$, then $X$ has a Hamilton cycle by Theorem 1.1 in [17] and Theorem 2.5 in [25], respectively. From now on, suppose we are in Case (1) and $p\geq 5$. Let $N\neq 1$ be a maximal intransitive normal subgroup of $G$ on $V(X)$, that is, every normal subgroup of $G$ containing $N$ is transitive on $V(X)$ except for $N$. Let $\mathcal{B}$ be the set of $N$-blocks. Then $G/N$ acts quasiprimitively on $\mathcal{B}$.
If $|B|=6$ or $p$ (the case $p=5$ was included), then $X$ has a Hamilton cycle by Lemmas 3.5 and 3.4 in [19], respectively.
The remaining cases that $|B|=3p$, $|B|\in\{2,3\}$ and $|B|=2p$ will be dealt with in the following three subsections and it will be proved that $X$ contains a Hamilton cycle, except for the triangle-replaced graph of the Petersen graph, see Theorems 3.1,  3.10,  3.11 and 3.12. Thus, Theorem 1.1 is proved. $\Box$
3.1 $|B|=3p$
Theorem 3.1
If $|B|=3p$, then $X$ has a Hamilton cycle.
Proof   Under the hypothesis, one may set =$\mathcal{B}=\{B_{0},B_{1}\}$ and $N=G_{B_{i}}$ (the setwise block stabilizer), where $i\in\mathbb{Z}_{2}$. Then $G=N\langle\tau\rangle$, for some $\tau$ of order $2^{e}$, which swaps the two blocks. Then we have four cases: (1) $N$ is unfaithful on some $N$-block;
(2) $N$ is faithful on both $N$-blocks and is primitive on some $N$-block;
(3) $N$ is faithful, imprimitive on both $N$-blocks but is not quasiprimitive on some $N$-block; and
(4) $N$ is faithful, imprimitive and quasiprimitive on both $N$-blocks. Case (1) will be dealt with in Lemma 3.2 and Remark 3.3; Case (2) will be dealt with in Lemma 3.4; Case (3) will be dealt with in Lemmas 3.5, 3.7 and Remark 3.6; and Case (4) will be dealt with in Lemma 3.8; separately. In all cases, $X$ will have a Hamilton cycle so that Theorem 3.1 holds. $\Box$
Lemma 3.2
Suppose that $N$ is unfaithful on some $N$-block and $p\neq 5$. Then $X$ has a Hamilton cycle.
Proof   Suppose that $N$ is unfaithful on an $N$-block, say $B_{0}$. Let $T_{0}=N_{(B_{0})}$ (the pointwise stabilizer) and $T=T_{0}\times T_{0}^{\tau}$. Then $T\lhd G$. If $T$ is transitive on both $B_{0}$ and $B_{1}$, then $X[B_{0},B_{1}]\cong K_{3p,3p}$ and so it has a Hamilton cycle. So suppose that $T$ induces blocks of length $r$, where $r\in\{3,p\}$. Let $\mathcal{B}^{\prime}$ be the set of $T$-blocks. By Proposition 2.6, the $r$-block graph $X_{\mathcal{B}^{\prime}}$ of order $\frac{6p}{r}$ (a product of two primes) has a Hamilton cycle. Take a Hamilton cycle $C$ of $X_{\mathcal{B}^{\prime}}$. Recall that $T$ is unfaithful on some $T$-block. Then there exist two adjacent $T$-blocks $B_{i}^{\prime}$ and $B_{j}^{\prime}$ in $C$ such that $d(B_{i}^{\prime},B_{j}^{\prime})=r$ in $X$. By Proposition 2.4, $C$ can be lifted to a Hamilton cycle of $X$. $\Box$
Remark 3.3
Suppose that $p=5$. Then the arguments in Lemma 3.2 are still true, except for $r=3$ and the $3$-block graph $X_{\mathcal{B}^{\prime}}$ is isomorphic to the Petersen graph. For this case, it will be shown in Lemma 3.7 that $X$ has a Hamilton cycle too.
Lemma 3.4
Suppose that $N$ is faithful on both $N$-blocks and is primitive on some $N$-block. Then $X$ has a Hamilton cycle.
Proof   Under the hypothesis, we may assume that $N$ is primitive on $B_{0}$. Take $u_{0}\in B_{0}$. Since $N_{u_{0}}$, the stabilizer of $u_{0}$ in $N$, is maximal in $N$ if and only if $N_{u_{0}^{\tau}}$ is maximal in $N$, we know that $N$ is primitive on $B_{1}$. Set $K=\hbox{\rm Soc}(N)$. Then $K~{}\hbox{\rm char}~{}N\lhd G$, $K\lhd G$, and so $K$ is transitive on both $B_{0}$ and $B_{1}$. Let $G_{0}=K\langle\tau\rangle$ and $K_{0}=K\langle\tau^{2}\rangle$. Then $G_{0}$ is transitive on $V(X)$ and $G_{0}/K_{0}\cong\mathbb{Z}_{2}$. By Proposition 2.14, $K\cong L$, where $L\in\{A_{3p}$, $A_{6}$, $A_{7}$, $\hbox{\rm PSL}(2,7)$, $\hbox{\rm PSL}(2,19),\hbox{\rm PSL}(d,q)\}$; and $K$ contains a metacirculant subgroup $T\cong\mathbb{Z}_{p}\rtimes\mathbb{Z}_{3^{i}}$ or $\mathbb{Z}_{3p}$, which is transitive on some $N$-block, say $B_{0}$, except for $K\cong A_{6}$ of degree $15$.
Firstly, suppose that $K\cong A_{3p}$, which is of degree $3p$. Then $K_{u_{0}}\cong A_{3p-1}$ and $K$ has two suborbits with length 1 and $13p-$ 1 on $B_{i}$, where $i\in\mathbb{Z}_{2}$. Therefore, $d(X)$, the valency of $X$, is at least $3p-1\geq\frac{6p}{3}$. By Proposition 2.1, $X$ has a Hamilton cycle.
Secondly, suppose that $K\cong A_{6}$, which is of degree $15$. Since $K$ is a nonabelian simple group, Aut ($\hbox{\rm Aut\,}(K)\cong P\Gamma L(2,9)$, Aut ($\hbox{\rm Aut\,}(K)/K\cong\mathbb{Z}_{2}^{2}$ and $G_{0}/K$ is cyclic, we get that $G_{0}=K\times\langle\tau\rangle\cong A_{6}\times\mathbb{Z}_{2}$, $G_{0}\cong S_{6}$ or $G_{0}\cong\hbox{\rm PGL}(2,9)$. For the first case, $\langle\tau\rangle$ induces blocks of length $2$. Since the $2$-block graph induced by $\langle\tau\rangle$ is a connected vertex-transitive graph of order $15$ with a Hamilton cycle and $G_{0}/\langle\tau\rangle$ is primitive on the set of $\langle\tau\rangle$-blocks, by Lemma 2.10, $X$ has a Hamilton cycle. For the last two cases, $X$ is isomorphic to a connected orbital graph $X_{0}=X(G_{0},H,D)$ of order $30$, where $G_{0}\cong S_{6}$ or $\hbox{\rm PGL}(2,9)$, and $H\cong S_{4}$. Checking all cases with Magma, we get that $X$ has a Hamilton cycle.
Finally, suppose that $K\cong L$, where $L\in\{A_{7},\hbox{\rm PSL}(2,7),\hbox{\rm PSL}(2,19),\hbox{\rm PSL}(d,q)\}$. Let $P\in\hbox{\rm Syl}_{p}(T)$ (also, $P\in\hbox{\rm Syl}_{p}(G_{0})$), and let $H=(K_{0})_{u_{0}}=(G_{0})_{u_{0}}$. Then $K_{0}=HT$. Recall that $\tau$ swaps $B_{0}$ and $B_{1}$. Then $B_{i}=u_{0}^{\tau^{-i}K_{0}}=u_{0}^{K_{0}\tau^{-i}}=\{u_{0}^{k\tau^{-i}}:k\in K_{0}\}$, where $i\in\mathbb{Z}_{2}$. We claim that $N_{K_{0}}(P)$ is transitive on $B_{1}$. In fact, there exists $g\in K_{0}$ such that $P^{\tau}=P^{g}$, that is $g\tau^{-1}\in N_{G_{0}}(P)$. Set $u_{1}=u_{0}^{g\tau^{-1}}$. Then $u_{1}\in B_{1}$. Since $G_{0}=K_{0}\cup K_{0}\tau^{-1}$ and $K_{0}\tau^{-1}=HT\tau^{-1}=HTg\tau^{-1}=Hg\tau^{-1}T^{g\tau^{-1}}$, we get that $B_{1}=u_{0}^{K_{0}\tau^{-1}}=u_{0}^{Hg\tau^{-1}T^{g\tau^{-1}}}=u_{1}^{T^{g\tau^{-1}}}$ and so $T^{g\tau^{-1}}$ is transitive on $B_{1}$ with a normal subgroup $P$. Since $TT^{g\tau^{-1}}\subset N_{K_{0}}(P)$, we get that $N_{K_{0}}(P)$ is transitive on both $B_{0}$ and $B_{1}$. Since $N_{G_{0}}(P)K_{0}=G_{0}$, we get that $N_{G_{0}}(P)\cap K_{0}\tau\neq\emptyset$ and so there exists $g_{1}\in N_{G_{0}}(P)$ such that $B_{0}^{g_{1}}=B_{1}$. Moreover, $N_{K_{0}}(P)\leq N_{G_{0}}(P)$, which implies that $N_{G_{0}}(P)$ is transitive on $V(X)$ containing a normal Sylow-$p$ subgroup $P$ which induces blocks of length $p$. By Proposition 2.9, $X$ has a Hamilton cycle. $\Box$
Lemma 3.5
Suppose that $p\neq 5$ and $N$ is faithful, imprimitive on both $N$-blocks but is not quasiprimitive on some $N$-block. Then $X$ has a Hamilton cycle.
Proof   Under the hypothesis, let $K\neq 1$ be a maximal intransitive normal subgroup of $N$ on an $N$-block, say $B_{0}$. Then $K$ induces blocks of length $p$ or 3 on $B_{0}$.
(1) Suppose that $K$ induces blocks of length $p$ on $B_{0}$. Then $N/K\lessapprox S_{3}$ and $K\lessapprox S_{p}^{3}$. Recall that $\tau$ swaps $B_{0}$ and $B_{1}$. Then $K^{\tau}$ induces blocks of length $p$ on $B_{1}$. Let $P\in\hbox{\rm Syl}_{p}(K)=\hbox{\rm Syl}_{p}(K^{\tau})=\hbox{\rm Syl}_{p}(G)$. Since $N$ is faithful on both $N$-blocks, $N_{N}(P)K=N$ and $N_{N}(P)K^{\tau}=N$, we know that $N_{N}(P)$ is transitive on both $B_{0}$ and $B_{1}$. Since $N_{G}(P)N=G$ and $N_{N}(P)\leq N_{G}(P)$, using the same argument as in Lemma 3.4, we know that $N_{G}(P)$ is transitive on $V(X)$ and contains a normal Sylow-$p$ subgroup $P$ which induces blocks of length $p$. By Proposition 2.9, $X$ has a Hamilton cycle.
(2) Suppose that $K$ induces blocks of length $3$ on $B_{0}$. Then $K\lessapprox S_{3}^{p}$ and $KK^{\tau}/K\lhd N/K\lessapprox S_{p}$. Since $p\nmid|KK^{\tau}/K|$, we get that $K=K^{\tau}$ and so $K\lhd G$. Let ${\mathcal{B}^{\prime}}$ be the set of $K$-blocks. Then the 3-block graph $X_{\mathcal{B}^{\prime}}$ is a connected vertex-transitive graph of order $2p$, where $p\geq 7$. By Proposition 2.6, $X_{\mathcal{B}^{\prime}}$ has a Hamilton cycle. Firstly, suppose that $K$ is unfaithful on some $K$-block of length $3$. Then using the arguments as in Lemma 3.2, every Hamilton cycle in $X_{\mathcal{B}^{\prime}}$ can be lifted to a Hamilton cycle of $X$. Secondly, suppose that $K$ is faithful on every $K$-block. Then $K\lessapprox S_{3}$. Take a subgroup $M/K$ of order $p$ in $N/K$. Then $M=K\times P$ (as $p\textgreater|S_{3}|$), for some $P\in\hbox{\rm Syl}_{p}(G)=\hbox{\rm Syl}_{p}(N)$, and $M$ is transitive on both $B_{0}$ and $B_{1}$. Moreover, $P^{\tau}=P^{g}$ for some $g\in N$, which implies that $M^{\tau g^{-1}}=M$. Thus, $G_{0}:=(K\times P)\langle\tau g^{-1}\rangle$ is transitive on $V(X)$ with a normal subgroup $P$. By Proposition 2.9 again, $X$ has a Hamilton cycle. $\Box$
Remark 3.6
The arguments in Lemma 3.5 are still true provided either $p=5$ and $K$ induces blocks of length $5$; or $K$ induces blocks of length $3$ but $X_{\mathcal{B}^{\prime}}$ is not isomorphic to the Petersen graph.
Lemma 3.7
Suppose that $p=5$ and $N$ is faithful, imprimitive on both $N$-blocks but is not quasiprimitive on some $N$-block. Then $X$ has a Hamilton cycle.
Proof   By Remarks 3.3 and 3.6, we only need to consider the case that $G$ has a normal subgroup which induces blocks of length $3$ and the $3$-block graph is isomorphic to the Petersen graph.
Let =${\mathcal{B}^{\prime}}=\{B_{ij}:i\in\mathbb{Z}_{2},j\in\mathbb{Z}_{5}\}$ be the set of blocks of length 3 and $T$ be the kernel of $G$ on ${\mathcal{B}^{\prime}}$, where $B_{ij}\subset B_{i}$. Then $T\lhd N$, $T\lessapprox S_{3}^{5}\cong\mathbb{Z}_{3}^{5}\rtimes\mathbb{Z}_{2}^{5}$ and $N/T\lessapprox S_{5}$. Let $\mathcal{B}_{i}^{\prime}=\{B_{ij}:j\in\mathbb{Z}_{5}\}$, where $i\in\mathbb{Z}_{2}$. Since $X_{\mathcal{B}^{\prime}}$ is isomorphic to the Petersen graph, we know that $X_{\mathcal{B}_{i}^{\prime}}\cong C_{5}$, the cyclic graph of order $5$, is edge-transitive and $G/T\lessapprox S_{5}$. Moreover, $G/T$ is a non-quasiprimitive group of degree $10$ containing a nontrivial normal subgroup $N/T$, we get that $G/T\cong\hbox{\rm AGL}(1,5)$ and $N/T\cong\mathbb{Z}_{5}\rtimes\mathbb{Z}_{2}$. Set $P\in\hbox{\rm Syl}_{3}(T)=\hbox{\rm Syl}_{3}(G)$ and $U\in\hbox{\rm Syl}_{3}(S_{3}^{5})$. Then $U\lhd S_{3}^{5}$. Since $N_{G}(P)T=G$, we know that $N_{G}(P)$ is transitive on $V(X)$ with a normal Sylow 3-subgroup $P$ as in Lemma 3.5. Since $\langle P,b\rangle\lessapprox U$, for any $b\in T$ of order $3$, we know that $\langle P,b\rangle$ is also a $3$-subgroup of $T$ and so $P=\langle P,b\rangle\lhd T\lhd N_{G}(P)$. Since $N_{G}(P)/T\leq G/T$ and $X_{\mathcal{B}^{\prime}}$ is not a Cayley graph, we know that $N_{G}(P)=G$. Clearly, $G/P$ acts naturally on $\mathcal{B}^{\prime}$ with the kernel $T/P$. Let $Q\in\hbox{\rm Syl}_{5}(N)=\hbox{\rm Syl}_{5}(G)$. Set $\overline{Q}=QP/P$, $\overline{N}=N/P$ and $\overline{G}=G/P$. Then $\overline{Q}\in\hbox{\rm Syl}_{5}(\overline{N})=\hbox{\rm Syl}_{5}(\overline{G})$ and $N_{\overline{G}}(\overline{Q})\overline{N}=\overline{G}$. Thus, $N_{\overline{G}}(\overline{Q})$ is transitive on $\mathcal{B}^{\prime}$. Let $G_{0}$ be the pre-image of $N_{\overline{G}}(\overline{Q})$ in $G$ containing $P$. Then $G_{0}$ is transitive on $V(X)$ and has a normal subgroup $P\rtimes Q$. Moreover, there exists $d\in G_{0}$ such that $B_{0}^{d}=B_{1}$ for some $d$ of order $2^{e}$, where $e\in\{2,3\}$. Set $D=\langle d\rangle$ and $G_{1}=PQD=(P\rtimes Q)\rtimes D\cong(P\rtimes\mathbb{Z}_{5})\rtimes\mathbb{Z}_{2^{e}}\cong P\rtimes(\mathbb{Z}_{5}\rtimes\mathbb{Z}_{2^{e}})$. Then $G_{1}$ is transitive on $V(X)$, $\mathcal{B}^{\prime}$ is also the set of $P$-blocks and $\mathcal{B}$ is also the set of $PQ$-blocks.
Suppose that $P$ is faithful on some $P$-block. Then $P\cong\mathbb{Z}_{3}$ and $G_{1}\cong\mathbb{Z}_{5}\rtimes(\mathbb{Z}_{3}\rtimes\mathbb{Z}_{2^{e}})$. By Proposition 2.9, $X$ has a Hamilton cycle.
Suppose that $P$ is unfaithful on some $P$-block. Then $P\cong\mathbb{Z}_{3}^{n}$, where $1~{}\textless~{}n\leq 5$. Without loss of generality, we may set $Q=\langle c\rangle\in\hbox{\rm Syl}_{5}(G_{1})$, where $c^{d}=c^{t}$ for some $t\in\mathbb{Z}_{5}\backslash\{0\}$. Let $[P,Q]$ be the commutator subgroup of $P$ and $Q$. Suppose that $[P,Q]=1$. Then $G_{1}=Q\rtimes(P\rtimes D)\cong\mathbb{Z}_{5}\rtimes(\mathbb{Z}_{3}^{n}\rtimes\mathbb{Z}_{2^{e}})$. By Proposition 2.9, $X$ has a Hamilton cycle. Set $P_{z}=P\cap Z(G_{1})$, where $Z(G_{1})$ is the center of $G_{1}$. Then $P_{z}\lhd G_{1}$. If $P_{z}\neq 1$, then $G_{1}$ has a transitive subgroup $P_{z}\times(Q\rtimes D)\cong\mathbb{Z}_{5}\rtimes(P_{z}\times\mathbb{Z}_{2^{e}})$ on $V(X)$. By Proposition 2.9, $X$ has a Hamilton cycle. The case that both $[P,Q]\neq 1$ and $P_{z}=1$ will be dealt with in the following.
In this case, we can regard $P$ as an $n$-dimensional linear space over $\mathbb{F}_{3}$, the finite field of order 3, and $c$ as an invertible matrix of an invertible linear transformation with respect to a basis of $P$, say $A$. Since $c^{5}=1$, we know that $g(x)=x^{5}-1=g_{1}(x)g_{2}(x)$ is an annihilating polynomial of $A$, where $g_{1}(x)=x-1$, $g_{2}(x)=\sum_{i=0}^{4}x^{i}$ and $g_{2}(x)$ is irreducible over $\mathbb{F}_{3}$. Let $f(x)$ be the characteristic polynomial of $A$. Then $g_{1}(x)\mid f(x)$ or $g_{2}(x)\mid f(x)$. Firstly, suppose that $n\in\{2,3\}$. Then $5\nmid|\hbox{\rm GL}(n,3)|$ and so $[P,Q]=1$, a contradiction. Secondly, suppose that $n=5$. Then 1 is an eigenvalue of $A$. Let $a\in P$ be an eigenvector of $A$ corresponding to the eigenvalue $1$. Then $a^{c}\in\langle a\rangle$ and so $ac=ca$. Set $a^{D}=\langle a^{x}\mid x\in D\rangle\leq P$. Since $a^{d^{i}c}=a^{c^{d^{-i}}d^{i}}=a^{d^{i}}$, we get that $[a^{D},Q]=1$ and $a^{D}\lhd G_{1}$. Set $G_{2}=(a^{D}\times Q)\rtimes D$. Then $G_{2}$ is transitive on $V(X)$ and $G_{2}=Q\rtimes(a^{D}\rtimes D)\cong\mathbb{Z}_{5}\rtimes(a^{D}\rtimes\mathbb{Z}_{2^{e}})$. By Proposition 2.9, $X$ has a Hamilton cycle. Finally, suppose that $n=4$. If $f(x)\neq g_{2}(x)$, using the same argument as above, then $P$ contains an eigenvector of $A$ corresponding to the eigenvalue $1$ and $X$ has a Hamilton cycle. Now we assume that $f(x)=g_{2}(x)$. Recall that $5\nmid|\hbox{\rm GL}(3,3)|$, and we know that the invariant subspace of the matrix $A$ is the trivial subspace of $P$ and so $P$ is faithful on $B_{0}$ and $B_{1}$. Thus, $P_{(B_{00})}\cong\mathbb{Z}_{3}^{3}$ is transitive on some $P$-block in $\mathcal{B}_{0}^{\prime}$ and so there exist two distinct $P$-blocks in $\mathcal{B}_{0}^{\prime}$, say $B_{0i}$ and $B_{0j}$, such that $X[B_{0i},B_{0j}]\cong K_{3,3}$, where $i,j\in\mathbb{Z}_{5}$. Recall that $X_{\mathcal{B}_{0}^{\prime}}\cong C_{5}$ is an edge-transitive graph, and we know that $X(B_{i})$, where $i\in\mathbb{Z}_{2}$, contains a subgraph isomorphic to $K_{3,3,3,3,3}$, a complete 5-partite graph of order 15. Without loss of generality, we may assume that $B_{00}^{d^{j}}=B_{0j}$ and $B_{0k}\sim B_{1k}$ in $X_{\mathcal{B}^{\prime}}$, where $j,k\in\mathbb{Z}_{5}$. By Lemma 2.3, we know that $P$ has a $(10,3)$-semiregular automorphism, say $\rho$. Since $X(B_{i})$ contains a subgraph isomorphic to $K_{3,3,3,3,3}$ and $X_{\mathcal{B}_{i}^{\prime}}\cong C_{5}$, where $i\in\mathbb{Z}_{2}$, we know that the restriction of $\rho$ on $B_{0k}\cup B_{1k}$, say $\rho_{k}$, can still be regarded as an automorphism of $X$, where $k\in\mathbb{Z}_{5}$. Let $P_{1}=\langle\rho_{k}\mid k\in\mathbb{Z}_{5}\rangle\cong\mathbb{Z}_{3}^{5}$. If $P_{1}^{d}\neq P_{1}$, recall that $B_{0}^{d}=B_{1}$, then there exist $z_{1}\in P_{1}$ and $z_{2}\in P_{1}^{d}$ such that $\langle z_{1}z_{2}\rangle$ acts trivially on $B_{0}$ and is transitive on some block in $\mathcal{B}_{1}^{\prime}$, which implies that $X[B_{0k},B_{1k}]\cong K_{3,3}$, where $k\in\mathbb{Z}_{5}$. Thus, Aut ($\hbox{\rm Aut\,}(X)$ has a transitive subgroup isomorphic to $\mathbb{Z}_{5}\rtimes(\mathbb{Z}_{3}\rtimes\mathbb{Z}_{4})$. By Proposition 2.9, $X$ has a Hamilton cycle. If $P_{1}^{c}\neq P_{1}$, using the same argument as above, then $X$ has a Hamilton cycle. If $P_{1}^{d}=P_{1}$ and $P_{1}^{c}=P_{1}$, then $G_{2}:=P_{1}\rtimes(Q\rtimes D)\cong(\mathbb{Z}_{3}^{5}\rtimes\mathbb{Z}_{5})\rtimes\mathbb{Z}_{2^{e}}$ is transitive on $V(X)$, coming back to the case that $n=5$. $\Box$
Lemma 3.8
Suppose that $N$ is faithful, imprimitive and quasiprimitive on both $N$-blocks. Then $X$ has a Hamilton cycle.
Proof   Let $K=\hbox{\rm Soc}(N)$. Under the hypothesis, by Proposition 2.12, we know that $M\cong\hbox{\rm PSL}(d,q)$ of degree $p=(q^{d}-1)/(q-1)$, where $d$ is a prime and $(d,p-1)=1$.
Suppose that $d\neq 2$. Take $v\in V(X)$. Remind that there exists a unique maximal subgroup of $K$, say $K_{0}$, such that $K_{v}\leq K_{0}$, $K$ has only one conjugacy class of subgroups isomorphic to $K_{0}$ and $K_{0}$ has only one conjugacy class of subgroups isomorphic to $K_{v}$. Thus, $K$ has only one conjugacy class of subgroups isomorphic to $K_{v}$. Let $B_{i,j}$ be the nontrivial block for $K$ on $B_{i}$, and let $\mathcal{B}_{i}^{\prime}=\{B_{i,j}\mid j\in\mathbb{Z}_{p}\}$, where $i\in\mathbb{Z}_{2}$ and $j\in\mathbb{Z}_{p}$. For any $u\in V(X)$, there exists $g\in K$ such that $K_{v}=K_{u}^{g}$. Let $P\in\hbox{\rm Syl}_{p}(N)$. Suppose that $N$ has a metacyclic subgroup which contains $P$ and is transitive on $B_{0}$. Then using the same argument as in Lemma 3.4, $N_{G}(P)$ is transitive on $V(X)$ containing a normal Sylow $p$-subgroup $P$. By Proposition 2.9, $X$ has a Hamilton cycle. Suppose that there exists no such transitive subgroup. Then it can be extracted from [32] and [33, Theorem 2.1] that $d(B_{i_{1},j_{1}},B_{i_{2},j_{2}})=3$ if $K_{B_{i_{1},j_{1}}}\neq K_{B_{i_{2},j_{2}}}$, where $i_{1},i_{2}\in\mathbb{Z}_{2}$, $j_{1},j_{2}\in\mathbb{Z}_{p}$. Without loss of generality, we may assume that $K_{B_{i_{1},j}}=K_{B_{i_{2},j}}$, where $j\in\mathbb{Z}_{p}$. Moreover, $X$ is connected implying that there exist $j\in\mathbb{Z}_{p}$ and $j\neq 0$ such that $B_{0,0}\sim B_{i,j}$, where $i\in\mathbb{Z}_{2}$. Since $K$ is 2-transitive on $\mathcal{B}_{i}^{\prime}$, where $i\in\mathbb{Z}_{2}$, we get that $d(X)\geq 3(p-1)\geq 2p$, by Proposition 2.1, $X$ has a Hamilton cycle.
Now, we may assume that $d=2$ and $K:=\hbox{\rm Soc}(N)=\hbox{\rm SL}(2,q)$. Since $K~{}\hbox{\rm char}~{}N\lhd G$, we know that $G_{0}:=K\langle\tau\rangle$ acts transitively on $V(X)$ as before. Let $\mathbb{F}_{q}^{*}=\langle\theta\rangle$, the multiplicative group of the finite field $\mathbb{F}_{q}$, and let $f(x)=x^{2}+\theta^{m}x+1$ be an irreducible polynomial over $\mathbb{F}_{q}$, for some $m\in\mathbb{Z}_{q-1}$. Set
$\begin{array}[]{ll}&\ell={\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right]},\,t=\left[\begin{array}[]{cc}\theta&0\\ 0&\theta^{-1}\end{array}\right],\,u=\left[\begin{array}[]{cc}1&1\\ 0&1\end{array}\right],\,s(a,b)=\left[\begin{array}[]{cc}a&b\\ b&a+b\theta^{m}\end{array}\right],\\ &\\ &S=\langle s(a,b)\bigm{|}a,b\in\mathbb{F}_{q},a^{2}+b^{2}+ab\theta^{m}=1\rangle.\end{array}$
Then ${\rm o}(t)=q-1,\,{\rm o}(u)=2,\,t^{\ell}=t^{-1}\,\,{\rm and}\,\,S\cong\mathbb{Z}_{q+1}.$ Remind that $K$ has only one conjugacy class of subgroups isomorphic to $S$.
Let $T={(G_{0})}_{\infty}=\langle u,t\rangle\cong\mathbb{Z}_{2}^{2^{s}}\rtimes\mathbb{Z}_{q-1}$, the point stabilizer of $G_{0}$ relative to $\infty$ in the projective line $\hbox{\rm PG}(1,q)$. Let $H=\langle u,t^{3}\rangle\leq T$ and $[K:H]$ be the set of right cosets of $K$ relative to $H$. Then $K$ has only one conjugacy class of subgroups isomorphic to $H$ and $H$ is the stabilizer of some vertex of $V(X)$ in $K$. Moreover, one may set $\tau\in N_{G_{0}}(T)$. We shall prove this case by the following three steps.
Step 1: Determination of suborbits of $K$.
Let $\alpha=H\in[K:H]$. Since the group $T$ has two orbits on $[K:T]$ (as $K$ is 2-transitive on $\hbox{\rm PG}(1,q)$) and $T=H\cup Ht\cup Ht^{2}$, we get that the group $K$ has totally three single point suborbits $\{\alpha^{t^{i}}\}$ and three suborbits $\alpha^{t^{i}\ell H}$ of length $q$, where $i\in\mathbb{Z}_{3}$. Since $t^{i}\ell$ is an involution, the suborbit $\alpha^{t^{i}\ell H}$ is self-paired for any $i\in\mathbb{Z}_{3}$.
Since $H\cap S\leq T\cap S=1$, the group $S$ acts semiregularly on $[K:H]$. Since the group $S$ acts regularly on $[K:T]$, by $T=\cup_{i=0}^{2}Ht^{i}$ again, we know that the group $S$ has three orbits, say $\alpha^{t^{i}S}$, on $[K:H]$, where $i\in\mathbb{Z}_{3}$, and all of them are of length $q+1$.
Step 2: Determine some orbital graphs arising from the action of $K$ on $[K:H]$.
Firstly, consider the orbital graph $Y_{i}=X(K,H,Ht^{i}\ell H)$ with $i\in\mathbb{Z}_{3}$, and we shall show $d(\alpha^{t^{j}S},\alpha^{t^{k}S})\geq 2$ in $Y_{i}$, where $j,k\in\mathbb{Z}_{3}$.
Remind that $N_{1,i}(\alpha)$, the neighborhood of $\alpha$ in $Y_{i}$, is $\{\alpha^{t^{i}\ell h}:h\in H\}=\{Ht^{i}\ell h:h\in H\}$, where $H=\langle u,t^{3}\rangle$. Then
$\begin{array}[]{lcl}N_{1,i}(\alpha^{t^{j}})&=&\{Ht^{i}\ell ht^{j}:h\in H\}=\{Ht^{i-j}\ell h_{1}:h_{1}\in H\}\\ &=&\{H{\left[\begin{array}[]{cc}0&\theta^{i-j}\\ \theta^{-(i-j)}&x\end{array}\right]}:x\in\mathbb{F}_{k}\}.\\ \end{array}$
Remind that $d(\alpha^{t^{j}S},\alpha^{t^{k}S})=|N_{1,i}(\alpha^{t^{j}})\cap\alpha^{t^{k}S}|$, which is the number of solutions $s(a,b)$ of the equation
$H{\left[\begin{array}[]{cc}0&\theta^{i-j}\\ \theta^{-(i-j)}&x\end{array}\right]}=Ht^{k}s(a,b),$ (3)
that is
${\left[\begin{array}[]{cc}\theta^{3r}&x_{1}\\ 0&\theta^{-3r}\end{array}\right]}{\left[\begin{array}[]{cc}0&\theta^{i-j}\\ \theta^{-(i-j)}&x\end{array}\right]}=t^{k}s(a,b),$
for some $r\in\mathbb{Z}_{(q-1)/3}$ and some $x_{1}\in\mathbb{Z}_{q}$, that is
$\left[\begin{array}[]{cc}\theta^{-(i-j)}x_{1}&xx_{1}+\theta^{3r+i-j}\\ \theta^{-(3r+i-j)}&\theta^{-3r}x\end{array}\right]=\left[\begin{array}[]{cc}a\theta^{k}&b\theta^{k}\\ b\theta^{-k}&(a+b\theta^{m})\theta^{-k}\end{array}\right].$
Then Eq(3) holds if and only if
[…]
that is
$\left\{\begin{array}[]{lcll}x_{1}&=&a\theta^{i-j+k},\,\hskip 56.9055pt&(i^{\prime})\\ xx_{1}+\theta^{3r+i-j}&=&b\theta^{k},&(ii)\\ b&=&\theta^{-(3r+i-j-k)},&(iii^{\prime})\\ x&=&(a+b\theta^{m})\theta^{3r-k},&(iv^{\prime})\\ a^{2}+ab\theta^{m}+b^{2}&=&1.&(v)\end{array}\right.$
Inserting $(i^{\prime})$, $(iii^{\prime})$ and $(iv^{\prime})$ to $(ii)$, we obtain
$a(a+\theta^{-(3r+i-j-k)+m})\theta^{3r+i-j}+\theta^{3r+i-j}=\theta^{-(3r+i-j-2k)}.\hskip 28.45274pt(ii^{\prime})$
Note that $(ii^{\prime})$ and $(iii^{\prime})$ may imply $(v)$. Set $y=\theta^{-r}$ and $c=\theta^{-(i-j-k)}$. Then the equation $(ii^{\prime})$ becomes
$a^{2}+c\theta^{m}ay^{3}+c^{2}y^{6}+1=0.$ (4)
Conversely, given any solution $(a,y)$ of Eq(4), $x_{1}$, $b$ and $x$ are uniquely determined by $(i^{\prime})$, $(iii^{\prime})$ and $(iv^{\prime})$, respectively; clearly, $(ii^{\prime})$ holds and so $(ii)$ holds; and finally $(v)$ holds following $(ii^{\prime})$ and $(iii^{\prime})$ hold.
In summary, $d(\alpha^{t^{j}S},\alpha^{t^{k}S})\geq 2$ if and only if Eq(4) has at least two solutions. Let $n$ be the number of solutions of Eq(4). In what follows we shall show $n\geq 2$.
Firstly, suppose that Eq(4) is reducible in some field $\mathbb{F}\geq\mathbb{F}_{q}$. Then there exist $g(y)$ and $h(y)\in\mathbb{F}[y]$ such that $(a+g(y))(a+h(y))=0$. That is $g(y)+h(y)=c\theta^{m}y^{3}$ and $g(y)h(y)=c^{2}y^{6}+1$. Therefore, one has to set $g(y)=u_{3}y^{3}+u_{2}y^{2}+u_{1}y+u_{0}$ and $h(y)=v_{3}y^{3}+v_{2}y^{2}+v_{1}y+v_{0}$, where $u_{3}+v_{3}=c\theta^{m}$ and $u_{i}=v_{i}$ for $i\in\mathbb{Z}_{3}$. By $g(y)h(y)=c^{2}y^{6}+1$, we get
$\begin{array}[]{lcl}c^{2}y^{6}+1&=&g(y)h(y)=(u_{3}y^{3}+u_{2}y^{2}+u_{1}y+u_{0})(v_{3}y^{3}+u_{2}y^{2}+u_{1}y+u_{0})\\ &=&u_{3}v_{3}y^{6}+(u_{3}+v_{3})y^{3}(u_{2}y^{2}+u_{1}y+u_{0})+(u_{2}y^{2}+u_{1}y+u_{0})^{2}\\ &=&u_{3}v_{3}y^{6}+c\theta^{m}(u_{2}y^{5}+u_{1}y^{4}+u_{0}y^{3})+u_{2}^{2}y^{4}+u_{1}^{2}y^{2}+u_{0}^{2}\\ &=&u_{3}v_{3}y^{6}+c\theta^{m}u_{2}y^{5}+(c\theta^{m}u_{1}+u_{2}^{2})y^{4}+c\theta^{m}u_{0}y^{3}+u_{1}^{2}y^{2}+u_{0}^{2},\end{array}$
which implies that $c\theta^{m}u_{0}=0$ and $u_{0}^{2}=1$, a contradiction.
Secondly, suppose that Eq(4) is absolutely irreducible. Then it follows from Proposition 2.18 that
$n\geq q-20q^{\frac{1}{2}}-36.$ (5)
Suppose $q\geq 484$. Then $q-20q^{\frac{1}{2}}-36\geq 2$. Suppose $q\leq 483$. Since $q=2^{2^{s}}$, $q+1=p$ and $3\bigm{|}q-1$, we get $q=16$ or 256. Checking the two cases by Magma, Eq(4) has at least two solutions for any $c$, over $\mathbb{F}_{16}$ or $\mathbb{F}_{256}$.
Step 3: Graph $X$ has a Hamilton cycle.
Let $X$ be a connected vertex-transitive graph arising from $G_{0}$ acting on $[G_{0}:H_{0}]$, where $H_{0}=H\langle\tau^{2}\rangle$ is the stabilizer of some vertex of $V(X)$ in $G_{0}$ and $H_{0}\cap K=H$. Note that the group $K$ has two orbits on $[G_{0}:H_{0}]$ and the actions of $K$ on these two orbits are equivalent. Set ${\alpha_{l}}^{K}$ and ${\alpha_{r}}^{K}$ be these two orbits, respectively, where $\alpha_{l}=H_{L}$, $\alpha_{r}=H_{R}$ and $H_{L}=H_{R}=H_{0}$. Then $X$ can be seen as two parts, where one part is ${\alpha_{l}}^{K}$ and the other is ${\alpha_{r}}^{K}$. Note that $G_{0}$ has 12 suborbits on $[G_{0}:H_{0}]$, say $\{{\alpha_{l}}^{t^{i}}\},\{{\alpha_{r}}^{t^{i}}\},{\alpha_{l}}^{t^{i}\ell H_{0}}$ and ${\alpha_{r}}^{t^{i}\ell H_{0}}$, where $i\in\mathbb{Z}_{3}$. Similarly, the group $S$ has 6 orbits on $[G_{0}:H_{0}]$, saying ${\alpha_{l}}^{t^{i}S}$ and ${\alpha_{r}}^{t^{i}S}$, where $i\in\mathbb{Z}_{3}$, and all of them are of length $q+1$. Let $\mathcal{S}$ be the set of orbits of $S$. Since $\langle H_{0},\tau,t\rangle=\langle T,\tau\rangle\textless G_{1}$, we get that $X_{0}=X(G_{0},H_{0},D_{0})$ is disconnected, where $D_{0}$ is the union of all orbits of length $1$. Hence, the quotient graph $X_{\mathcal{S}}$ is also contains an edge $\{{\alpha_{l}}^{S},{\alpha_{r}}^{t^{i}lS}\}$ or $\{{\alpha_{l}}^{S},{\alpha_{r}}^{t^{i}lS}\}$.
${\alpha_{l}}^{S}$${\alpha_{l}}^{tS}$${\alpha_{l}}^{t^{2}S}$${\alpha_{r}}^{S}$${\alpha_{r}}^{tS}$${\alpha_{r}}^{t^{2}S}$
Figure 1: The quotient graph $X_{\mathcal{S}}$ of the vertex-transitive graph $X_{i}\cong X(G_{0},H_{0},H_{0}g\tau t^{i}\ell H_{0})$ arising from the action of $G_{0}$ on $[G_{0}:H_{0}]$ with respect to the $(6,p)$-semiregular automorphism $\rho=s(0,1)$, where ${\alpha_{l}}^{g\tau}=\alpha_{r}$ for some $g\in K$.
Suppose that ${\alpha_{r}}^{t^{i}\ell H_{0}}\subset N_{1}(\alpha_{l})$ for some $i\in\mathbb{Z}_{3}$, where $N_{1}(\alpha_{l})$ is the neighborhood of $\alpha_{l}$. Moreover, ${\alpha_{r}}^{t^{i}\ell H}\subset N_{1}(\alpha_{l})$. Recall that suborbits of $K$ on ${\alpha_{l}}^{K}$ and ${\alpha_{r}}^{K}$, which is determined in Step 1. Using the arguments in Step 2, one can see $X_{\mathcal{S}}$ contains a subgraph isomorphic to $K_{3,3}$, which is given in Figure 1, and every edge is a double edge (in fact, $d({\alpha_{l}}^{t^{i}S},{\alpha_{r}}^{t^{j}S})\geq 2$). Hence, we get a Hamilton cycle containing a double edge in $X_{\mathcal{S}}$:
${\alpha_{l}}^{S},{\alpha_{r}}^{S},{\alpha_{l}}^{tS},{\alpha_{r}}^{tS},{\alpha_{l}}^{t^{2}S},{\alpha_{r}}^{t^{2}S},\alpha_{l}^{S},$
By Proposition 2.4, it can be lifted to a Hamilton cycle of $X$, as desired.
Suppose that ${\alpha_{l}}^{t^{j}\ell H_{0}}\subset N_{1}(\alpha_{l})$ for some $j\in\mathbb{Z}_{3}$. Then using the same argument as in the previous paragraph, one can see each part of $X$ contains a subgraph isomorphic to $Y_{j}$, where $j\in\mathbb{Z}_{3}$. Due to the connectivity of $X$, there must be an edge between ${\alpha_{l}}^{K}$ and ${\alpha_{r}}^{K}$. We only need to consider that $\alpha_{l}\sim\alpha_{r}^{t^{i}}$, where $i\in\mathbb{Z}_{3}$. We get that $X_{\mathcal{S}}$, given in Figure 2, contains a perfect matching. Hence, we get a Hamilton cycle containing a double edge $\{{\alpha_{l}}^{S},{\alpha_{l}}^{tS}\}$ in $X_{\mathcal{S}}$:
${\alpha_{l}}^{S},{\alpha_{l}}^{tS},{\alpha_{l}}^{t^{2}S},{\alpha_{r}}^{t^{2+i}S},{\alpha_{r}}^{t^{1+i}S},{\alpha_{r}}^{t^{i}S},{\alpha_{l}}^{S},$
and so $X$ has a Hamilton cycle.
${\alpha_{l}}^{S}$${\alpha_{l}}^{tS}$${\alpha_{l}}^{t^{2}S}$${\alpha_{r}}^{t^{i}S}$${\alpha_{r}}^{t^{1+i}S}$$\alpha_{r}^{t^{2+i}S}$
Figure 2: The quotient graph $X_{\mathcal{S}}$ of the vertex-transitive graph $X_{i,j}\cong X(G_{0},H_{0},D)$ arising from the action of $G_{0}$ on $[G_{0}:H_{0}]$ with respect to the $(6,p)$-semiregular automorphism $\rho=s(0,1)$, where $D=H_{0}t^{j}\ell H_{0}\cup H_{0}g\tau t^{i}H_{0}\cup H_{0}t^{-i}(g\tau)^{-1}H_{0}$ and ${\alpha_{l}}^{g\tau}=\alpha_{r}$ for some $g\in K$.
Remark 3.9
In fact, Lemma 3.8 is also true when $K\cong\hbox{\rm SL}(2,q)$ and $K$ is a quasiprimitive group of order $pr$, where $p=q+1$ is a Fermat prime and $r$ is a small prime such as $r=5$ or $7$.
3.2 $|B|\in\{2,3\}$
Theorem 3.10
Suppose that $|B|=2$. Then $X$ has a Hamilton cycle.
Proof   Under the hypothesis, the 2-block graph $X_{\mathcal{B}}$ is a connected vertex-transitive graph of order $3p$. By Proposition 2.6, $X_{\mathcal{B}}$ has a Hamilton cycle. Let Aut ($\hbox{\rm Aut\,}(X_{\mathcal{B}})$ be the normalizer of $N$ in Aut ($\hbox{\rm Aut\,}(X)$. Then $G\in\hbox{\rm Aut\,}(X_{\mathcal{B}})$ Suppose that $N$ is unfaithful on a block in $\mathcal{B}$. Then every Hamilton cycle $C$ in $X_{\mathcal{B}}$ can be lifted to a Hamilton cycle of $X$ as in Lemma 3.2. Therefore, we assume that $N$ is faithful on each $N$-block, that is $N\cong\mathbb{Z}_{2}$. Recall that $G/N$ is a quasiprimitive group of degree $3p$, by Propositions 2.14 and 2.15, we get that either Aut ($\hbox{\rm Aut\,}(X_{\mathcal{B}})$ contains a metacyclic transitive subgroup $T\cong\mathbb{Z}_{p}\rtimes\mathbb{Z}_{3}^{e}$, $\mathbb{Z}_{3p}$ on $\mathcal{B}$; or $\hbox{\rm Soc}(G/N)\cong L$, where $L\in\{A_{6},\hbox{\rm SL}(2,q)\}$ and $p=q+1$ is a Fermat prime. Moreover, in the form case, the pre-image of Aut ($\hbox{\rm Aut\,}(X_{\mathcal{B}})$ in Aut ($\hbox{\rm Aut\,}(X)$ is isomorphic to $N.\hbox{\rm Aut\,}(X_{\mathcal{B}})$.
For the metacyclic cases, Aut ($\hbox{\rm Aut\,}(X)$ has a transitive subgroup $G_{0}\cong N.T$. Then either $G_{0}\cong\mathbb{Z}_{2}\times(\mathbb{Z}_{p}\rtimes\mathbb{Z}_{3}^{e})=\mathbb{Z}_{p}\rtimes\mathbb{Z}_{2\cdot 3^{e}}$ or $G_{0}\cong\mathbb{Z}_{6p}$. In both cases, $G_{0}$ has a normal Sylow $p$-subgroup which induces blocks of length $p$. By Proposition 2.9, $X$ has a Hamilton cycle.
Suppose that $L:=\hbox{\rm Soc}(G/N)\cong A_{6}$, which is of degree $15$. Then $L$ is transitive on $\mathcal{B}$ and so $G$ has a transitive subgroup $G_{0}\cong N.L\cong\mathbb{Z}_{2}.A_{6}$. Set $B_{0}\in\mathcal{B}$ and $u_{0}\in B_{0}$. Since the Schur multiplier of $A_{6}$ is $\mathbb{Z}_{6}$, we get that $G_{0}\cong\mathbb{Z}_{2}\times A_{6}$ or $G_{0}\cong\hbox{\rm SL}(2,9)$. In the former case, $X_{\mathcal{B}}$ is a connected vertex-transitive graph of order $15$ with a Hamilton cycle, by Lemma 2.10, $X$ has a Hamilton cycle. In the latter case, let $H=(G_{0})_{u_{0}}$. Then $(G_{0})_{B_{0}}=NH=N\times H$ and $N\leq N\times H\leq G_{0}$. Since $N$ is abelian and $(|N|,|G_{0}:NH|)=(2,3p)=1$, by Gaschutz’s Theorem ([14, I. Theorems 17.4 and 17.5]), $N$ has a complement in $G_{0}$, that is $\hbox{\rm SL}(2,9)\cong G_{0}\cong\mathbb{Z}_{2}\times A_{6}$, a contradiction.
Suppose that $L:=\hbox{\rm Soc}(G/N)\cong\hbox{\rm SL}(2,q)$, where $p=q+1$ is a Fermat prime. Then $L$ is transitive on $\mathcal{B}$ and so $G$ has a transitive subgroup $G_{0}\cong N.L\cong\mathbb{Z}_{2}.\hbox{\rm SL}(2,q)$. Since the Schur multiplier of $\hbox{\rm SL}(2,q)$ is $1$ when $q\neq 4$ or $q\neq 9$, we get that $G_{0}\cong N\times L$. Now the normal subgroup of $G_{0}$ which is isomorphic to $L$ induces blocks of length $3p$, coming back to the case in Subsection 3.1. $\Box$
Theorem 3.11
If $|B|=3$, then $X$ has a Hamilton cycle.
Proof   Now $|{\cal B}|=2p$ and $G/N$ acts faithfully and quasiprimitively on =$\mathcal{B}=\{B_{i}:i\in\mathbb{Z}_{2p}\}$. Then $X_{\mathcal{B}}$ is a connected vertex-transitive graph of order $2p$. We divide the proof into three cases.
(1) $G/N$ acts primitively on $\cal{B}$ and $\hbox{\rm Soc}(G/N)\not\cong A_{5}$.
By Proposition 2.6, $X_{\mathcal{B}}$ has a Hamilton cycle. As before, we assume that $N$ is faithful on each $N$-block, that is $A_{3}\lessapprox N\lessapprox S_{3}$. Let $G_{0}/N=\hbox{\rm Soc}(G/N)$. By Proposition 2.13, we know that $G_{0}/N\cong T$, where $T\in\{M_{22},\hbox{\rm PSL}(2,q),A_{2p}\}$. Let $Q$ be the Sylow-$3$ subgroup of $N$ and $C=C_{G_{0}}(Q)$. Then $Q$ char $N\lhd G_{0}$, $Q\lhd G_{0}$ and so $C\lhd G_{0}$. From $(G_{0}/Q)/(C/Q)\cong G_{0}/C\lessapprox\hbox{\rm Aut\,}(\mathbb{Z}_{3})\cong\mathbb{Z}_{2}$ and $C/Q\cong CN/N\lhd G_{0}/N$, we get that $[G_{0}:C]\leq 2$, $CN=G_{0}$ and $C\cong Q.T$. Thus, $C$ is transitive on $V(X)$ and $C/N$ is primitive on $\mathcal{B}$.
Suppose that $T=\hbox{\rm PSL}(2,9)$. Then $X_{\mathcal{B}}$ is a connected vertex-transitive graph of order $10$ containing a Hamilton cycle. Using the same argument as in Theorem 3.10, we know that $C\cong\mathbb{Z}_{3}\times A_{6}$. By Lemma 2.10, $X$ has a Hamilton cycle.
Suppose that $T=A_{2p}$ or $\hbox{\rm PSL}(2,q)$, where $q+1=2p\geq 14$. Then $X_{\mathcal{B}}\cong K_{2p}$ (as $C/Q$ is 2-transitive on $\mathcal{B}$). Since the Schur multiplier of $T$ is $\mathbb{Z}_{2}$, we have $C\cong T\times\mathbb{Z}_{3}$. By Lemma 2.10, $X$ has a Hamilton cycle.
Suppose $T=M_{22}$. Then $C\cong\mathbb{Z}_{3}.M_{22}$. Set $B_{0}\in\mathcal{B}$ and $u_{0}\in B_{0}$. Then $C_{B_{0}}=QC_{u_{0}}=Q\times C_{u_{0}}$. Thus, $Q\leq Q\times C_{u_{0}}\leq C$. By Gaschutz’s Theorem, $Q$ has a complement in $C$, that is $C\cong\mathbb{Z}_{3}\times M_{22}$. Moreover, $X_{\mathcal{B}}\cong K_{22}$, the complete graph of order $22$, has a Hamilton cycle. By Lemma 2.10 again, $X$ has a Hamilton cycle.
(2) $G/N$ acts primitively on $\cal{B}$ and $\hbox{\rm Soc}(G/N)\cong A_{5}$. Suppose that $K/N:=\hbox{\rm Soc}(G/N)\cong A_{5}$. Then $K/N$ is transitive on both $V(X_{\mathcal{B}})$ and $E(X_{\mathcal{B}})$. Moreover, $K$, the pre-image of $K/N$ in $G$, is transitive on $V(X)$.
Suppose that $X_{\mathcal{B}}$ has a Hamilton cycle. As before, we assume that $N\lessapprox S_{3}$. Let $Q\in\hbox{\rm Syl}_{3}(N)$ and $C=C_{K}(Q)$. Since the Schur multiplier of $A_{5}$ is $\mathbb{Z}_{2}$, using the same argument as in the previous case, we know that $C\cong\mathbb{Z}_{3}\times A_{5}$ is transitive on $V(X)$. By Lemma 2.10, $X$ has a Hamilton cycle. The remaining case that $X_{\mathcal{B}}$ is isomorphic to the Petersen graph will be discussed in the following.
Firstly, suppose that $N$ is faithful on some $N$-block. Then it is faithful on every $N$-block and so $A_{3}\lessapprox N\lessapprox S_{3}$. Let $P=\langle c\rangle\in\hbox{\rm Syl}_{3}(N)$ and $G_{0}=C_{K}(P)$. Using the same argument as before, we know that $G_{0}\cong\mathbb{Z}_{3}\times A_{5}$ is transitive on $V(X)$. Let $H=(G_{0})_{u_{0}}$, where $u_{0}\in V(X)$. Since $X_{\mathcal{B}}$ is isomorphic to the Petersen graph, $G_{0}$ has $6$ suborbits on $[G_{0}:H]$ meeting the condition, saying $\{Hc^{i}\}$ of length $1$ and $Hac^{i}H$ of length $3$ for some $a\in G_{0}\backslash P$, where $i\in\mathbb{Z}_{3}$ and $H\cong S_{3}$. Thus, $X$ has a connected subgraph isomorphic to $X_{0}=X(G_{0},H,D)$, where $D=HcH\cup Hc^{-1}H\cup HaH$ or $D=HacH\cup Hac^{2}H$. Let $G_{1}=G_{0}.\mathbb{Z}_{2}\cong\mathbb{Z}_{3}\times S_{5}$, and let $H_{1}\cong\mathbb{Z}_{2}\times S_{3}$ be a subgroup of $G_{1}$ containing $H$. Let $D_{1}=H_{1}DH_{1}$, a union of some double cosets of $H_{1}$, and let $X_{1}\cong X(G_{1},H_{1},D_{1})$. Let $f$ be a bijection from $V(X_{0})$ to $V(X_{1})$ such that $f(Hg)=H_{1}g$, where $g\in G_{0}$. Then we find that $Hg_{1}\sim Hg_{2}$ in $X_{0}$ if and only if $H_{1}g_{1}\sim H_{1}g_{2}$ in $X_{1}$, where $g_{1},g_{2}\in G_{0}$, and so $f$ is also an isomorphism from $X_{0}$ to $X_{1}$. Therefore, $\mathbb{Z}_{3}\times S_{5}$ is isomorphic to a transitive subgroup of Aut ($\hbox{\rm Aut\,}(X_{0})$. Moreover, Aut ($\hbox{\rm Aut\,}(X_{0})$ has a transitive subgroup which is isomorphic to $\mathbb{Z}_{3}\times(\mathbb{Z}_{5}\rtimes\mathbb{Z}_{4})=\mathbb{Z}_{5}\rtimes\mathbb{Z}_{12}$. By Proposition 2.9, $X$ has a Hamilton cycle.
Secondly, suppose that $N$ is unfaithful on some $N$-block. As before, there exist two adjacent $N$-blocks $B_{i}$ and $B_{j}$ such that $X[B_{i},B_{j}]\cong K_{3,3}$, where $i,j\in\mathbb{Z}_{10}$ and $i\neq j$. Since $X_{\mathcal{B}}$ is an edge-transitive graph, we get that if $B_{s}\sim B_{t}$, then $X[B_{s},B_{t}]\cong K_{3,3}$ and so $X$ has a subgraph which is isomorphic to $X(G_{1},H_{1},D_{1})$ in the preceding paragraph, where $s\neq t$ and $H_{1}aH_{1}\cup H_{1}acH_{1}\cup H_{1}ac^{2}H_{1}\subset{D_{1}}$. Thus, $X$ has a Hamilton cycle.
(3) $G/N$ acts quasiprimitively but imprimitively on $\cal{B}$. Checking Proposition 2.15, we know that Aut ($\hbox{\rm Aut\,}(X_{\mathcal{B}})$ contains a subgroup $G_{0}/N\cong\mathbb{Z}_{p}\rtimes\mathbb{Z}_{2s}$, which is transitive on $\mathcal{B}$, and $G_{0}$ is transitive on $V(X)$. Without loss of generality, we may take $s=2^{k-1}$ for some $k$. As before, we assume that $N$ is faithful on each $N$-block, that is $N\lessapprox S_{3}$. Let $P\in\hbox{\rm Syl}_{p}(G_{0})$. Then $PN/N\lhd G_{0}/N$ and so $PN\lhd G_{0}$. Moreover, $PN$ induces blocks of length $3p$, which is the case in Subsection 3.1. $\Box$
3.3 $|B|=2p$
Theorem 3.12
Suppose $|B|=2p$. Then $X$ has a Hamilton cycle.
Proof   Under the hypothesis, one may set =$\mathcal{B}=\{B_{0},B_{1},B_{2}\}$, $N=G_{B_{i}}$ and $G/N\cong\mathbb{Z}_{3}$. Then $G=N\langle\tau\rangle$, for some $\tau$ of order $3^{e}$, which swaps the three blocks. As in Subsection 3.1, we shall deal with four cases in the following Lemmas 3.13-3.16, separately. In all cases, $X$ has a Hamilton cycle. $\Box$
Lemma 3.13
Suppose that $N$ is unfaithful on some $N$-block. Then $X$ has a Hamilton cycle.
Proof   Suppose that $N$ is unfaithful on an $N$-block, say $B_{0}$. Let $T_{0}=N_{(B_{0})}$ and $T_{i}=T_{0}^{\tau^{i}}$, where $i\in\mathbb{Z}_{3}$.
Firstly, suppose that $T_{0}$ is transitive on $B_{1}$. Then $X[B_{0},B_{1}]\cong K_{2p,2p}$. Since $\langle\tau\rangle$ is transitive on the three $N$-blocks, we know that $X_{\mathcal{B}}$ is an edge-transitive graph and so $X$ contains a subgraph isomorphic to $K_{2p,2p,2p}$. Thus, $X$ has a Hamilton cycle.
Secondly, suppose that $T_{0}$ induces blocks of length $r$ on $B_{1}$, where $r\in\{2,p\}$. Suppose $r=p$. Then $d(B_{i},B_{j})=d(B_{0},B_{1})\geq p$, where $i\neq j$ and $i,j\in\mathbb{Z}_{3}$. The valency of $X$ is at least $2p$. By Proposition 2.1, $X$ has a Hamilton cycle. The case that $T_{0}$ induces blocks of length $p$ on $B_{2}$ is similar to the above. Next, we consider the remaining case that $T_{0}$ does not induce blocks of length $p$ on both $B_{1}$ and $B_{2}$. Suppose that $r=2$. Then $T_{0}$ is a 2-group and so the normal 2-subgroup $T=T_{0}T_{1}T_{2}$ of $G$ induces blocks of length 2. Let ${\mathcal{B}^{\prime}}$ be the set of $T$-blocks. Then $X_{\mathcal{B}^{\prime}}$ is a connected vertex-transitive graph of order $3p$ containing a Hamilton cycle. Since $T$ is unfaithful on some $T$-block, using the same argument as in Lemma 3.2, every Hamilton cycle of $X_{\mathcal{B}^{\prime}}$ can be lifted to a Hamilton cycle of $X$. $\Box$
Lemma 3.14
Suppose that $N$ is faithful on all $N$-blocks and is primitive on some $N$-block. Then $X$ has a Hamilton cycle.
Proof   Suppose that $N$ is primitive on an $N$-block, say $B_{0}$. Then it is primitive on the other two $N$-blocks as in Lemma 3.4. Set $K=\hbox{\rm Soc}(N)$. Then $K$ char $N\lhd G$ and so $K\lhd G$. Let $G_{1}=K\langle\tau\rangle$. Then $G_{1}$ is transitive on $V(X)$ as before. By Proposition 2.13, $K\cong T$, where $T\in\{A_{2p}$, $\hbox{\rm PSL}(2,q)$, $M_{22}$, $A_{5}\}$. Take $u_{0}\in B_{0}$.
Firstly, suppose $K\cong A_{2p}$. Then $K_{u_{0}}\cong A_{2p-1}$. Since $K$ has 2 suborbits of length 1, $12p-$ 1, respectively, on each $N$-block, we get that either $d(X)\geq(2p-1)+2\geq\frac{6p}{3}$ when $d(B_{i},B_{j})=1$; or $d(X)\geq 2(2p-1)\geq\frac{6p}{3}$ when $d(B_{i},B_{j})\geq 2$, where $i\neq j$. By Proposition 2.1, $X$ has a Hamilton cycle.
Secondly, suppose $K\cong\hbox{\rm PSL}(2,q)$, where $q=p_{0}^{t}=2p-1$. Then $K_{u_{0}}\cong\mathbb{Z}_{p_{0}}^{t}\rtimes\mathbb{Z}_{(p_{0}^{t}-1)/2}$, where $q+1=2p$. Since $K$ has 2 suborbits of length 1 and $q$, respectively, on each $N$-block, we get that either $d(X)\geq q+2\geq\frac{6p}{3}$ when $d(B_{i},B_{j})=1$; or $d(X)\geq 2q\geq\frac{6p}{3}$ when $d(B_{i},B_{j})\geq 2$. By Proposition 2.1, $X$ has a Hamilton cycle.
Thirdly, suppose $K\cong M_{22}$. Then $K_{u_{0}}\cong M_{21}$ and $|V(X)|=66$. Since $K$ has 2 suborbits of length 1, 21, respectively, on each $N$-block, we get $d(X)\geq 23$. By Proposition 2.1, $X$ has a Hamilton cycle.
Finally, suppose $K\cong A_{5}$. Since Aut ($\hbox{\rm Aut\,}(A_{5})=S_{5}\cong A_{5}\rtimes\mathbb{Z}_{2}$, $A_{5}$ is a nonabelian simple group and $G_{0}/K$ is cyclic, we get that $G_{0}=K\times\langle\tau\rangle\cong A_{5}\times\mathbb{Z}_{3}$. By Theorem 3.11, $X$ has a Hamilton cycle. $\Box$
Lemma 3.15
Suppose that $N$ is faithful, imprimitive on all $N$-blocks but is not quasiprimitive on some $N$-block. Then $X$ has a Hamilton cycle.
Proof   Under the hypothesis, let $K\neq 1$ be a maximal intransitive normal subgroup of $N$ on an $N$-block, say $B_{0}$. Then $K$ induces blocks of length $p$ or $2$ on $B_{0}$.
(1) Suppose that $K$ induces blocks of length $p$ on $B_{0}$. Then $N/K\cong\mathbb{Z}_{2}$ and $K\lessapprox S_{p}^{2}$. Let $P\in\hbox{\rm Syl}_{p}(K)=\hbox{\rm Syl}_{p}(N)$. Using the same argument as in Lemma 3.5, we know that $N_{G}(P)$ is transitive on $V(X)$, whose normal subgroup $P$ induces blocks of length $p$. By Proposition 2.9, $X$ has a Hamilton cycle.
(2) Suppose that $K$ induces blocks of length $2$ on $B_{0}$. Then $K\lessapprox\mathbb{Z}_{2}^{p}$ and $N/K\lessapprox S_{p}$. Let $\Omega=\hbox{\rm Syl}_{2}(N)$ and $Q=\cap_{A\in\Omega}A$. Then $K\leq Q~{}\hbox{\rm char}~{}N\lhd G$ and so $Q\lhd G$. Let ${\mathcal{B}^{\prime}}$ be the set of $Q$-blocks. The $2$-block graph $X_{\mathcal{B}^{\prime}}$ is a connected vertex-transitive graph of order $3p$ and so it contains a Hamilton cycle. If $Q$ acts unfaithfully on $\mathcal{B}^{\prime}$, then as before every Hamilton cycle in $X_{\mathcal{B}^{\prime}}$ can be lifted to a Hamilton cycle of $X$. So we assume that $Q\cong\mathbb{Z}_{2}$. Then $K=Q\lhd G$. Take $P\in\hbox{\rm Syl}_{p}(G)=\hbox{\rm Syl}_{p}(N)$. Then $P^{\tau}=P^{g}$ for some $g\in N$, that is $P^{\tau g^{-1}}=P$. Thus, $S:=(K\times P)\langle\tau g^{-1}\rangle$ is transitive on $V(X)$ and contains a normal Sylow $p$-subgroup $P$. By Proposition 2.9 again, $X$ has a Hamilton cycle. $\Box$
Lemma 3.16
Suppose that $N$ is faithful, imprimitive and quasiprimitive on all $N$-blocks. Then $X$ has a Hamilton cycle.
Proof   Let $K=\hbox{\rm Soc}(N)$. Then by Proposition 2.12, $K\cong\hbox{\rm PSL}(n,q)$, where $n$ is an odd prime and $(n,q-1)=1$. Let $P\in\hbox{\rm Syl}_{p}(N)$. Suppose that Suppose that $N$ has a metacyclic subgroup which contains $P$ and is transitive on $B_{0}$, then with the same argument as Lemma 3.4, $X$ has a Hamilton cycle. Suppose that there exists no such subgroup, with the same argument as Lemma 3.8, we get that either $d(X)\geq 2(2p-2)\geq 2p$ if $d(B_{0})\leq 1$; or $d(X)\geq 2p-2+2=2p$ if $d(B_{0})\geq 2$. By Proposition 2.1, in both two cases, $X$ has a Hamilton cycle. $\Box$